If x and y vary inversely, write an equation in simplest form that relates x and y when x=45 and y=3/5.
Answer by stanbon(57307)

x and y vary inversely means xy = k for some value of k.
You need to find that k-value.
Since y=3/5 when x = 45 you have this equation:
(45)(3/5) = k
k = 27
Therefore the equation you want is xy = 27
Cheers,
Stan H.